log(5)625+log(9)(x-3)=6 need help please
Answer by fractalier(1829)

Okay, from
log(5)625 + log(9)(x-3) = 6
let us evaluate the left hand log first...
we ask, "What exponent must we raise 5 to to get to 625?" Well, 4. So we have
4 + log(9)(x-3) = 6
log(9)(x-3) = 2
Now convert this to an exponential form...
9^2 = x - 3
81 = x - 3
x = 84